# Enhanced collective vibrations in granular materials

### v1

@@ I Introduction
The physics of the jamming transition provides a useful framework for understanding the complex dynamics of dense granular materials. Random packings of athermal, frictionless particles serve as the central model in this area of research  [19]. This model achieves mechanical stability at a specific density, known as the jamming transition [20, 21, 19, 22, 23, 24, 25, 26]. The jamming transition represents a form of the critical phenomenon, where various physical properties exhibit power laws in relation to the proximity to the critical density [27, 19, 23]. Among these, the vibrational dynamics of random packings have been extensively studied. The vibrational modes of the packings are derived from the diagonalization of the dynamical matrix, which depends solely on the mass, inter-particle interactions, and the configurations of the constituent particles. In crystals, these vibrational modes manifest as plane waves, and the vibrational density of states adhere to the notable Debye law $D(\omega)\propto\omega^{d-1}$, where $d$ denotes the spatial dimension. In contrast, many vibrational modes of the random packings are spatially disordered, contributing to an excess density of states relative to the Debye law. $D(\omega)$ exhibits a plateau down to a characteristic frequency that reveals critical power-law behavior near the jamming transition. Furthermore, the spatially localized vibrational modes, referred to as quasi-localized modes, are known to emerge in the lowest frequency regime [28, 29, 30]. Importantly, the linear rheological properties of the packings are dictated by these vibrational modes [31], highlighting the significance of understanding their vibrational properties.
@@ II.2 Linearization of the equation of motion
We consider the model at the volume fraction of $\phi\simeq 0.74$ which is higher than the jamming density. We focus on a mechanically stable configuration, or in other words, an inherent structure of the model. The position of particle $i$ in the inherent structure is denoted by ${\bm{R}}_{i}=(R_{ix},R_{iy},R_{iz})^{\top}$. Hereafter, we introduce a compact notation of a $3N$-dimensional vector like ${\bm{a}}=(\bm{a}_{1}^{\top},\bm{a}_{2}^{\top},\cdots,\bm{a}_{N}^{\top})^{\top}$. For example, the positions of particles in the inherent structure are denoted by ${\bm{R}}=(\bm{R}_{1}^{\top},\bm{R}_{2}^{\top},\cdots,\bm{R}_{N}^{\top})^{\top}$.
When the random force is sufficiently weak, the model behaves as a solid and the particles move only around their positions in the inherent structure. Hence, we linearize the equation of motion Eq. (1) in terms of the displacements of the particles, ${\bm{u}}\equiv{\bm{r}}-{\bm{R}}$, to obtain:
@@ II.4 Simulation setup
We first obtained the inherent structure of the model in the same way as Ref. [28]. The rattler particles which have less than $4$ contacting particles were removed. Starting from the inherent structures, we performed the molecular dynamics simulations of the linearized equation of motion Eq. (7). We performed sufficiently long simulations to obtain the steady state of the system, and then performed product runs where we calculated the physical quantities defined in Sec. II.3. To calculate the kinetic energies introduced by Eq. (14), we diagonalized the hessian matrix $\mathcal{H}$ to obtain the eigenvectors and the eigenfrequencies. We set the dissipation and random force parameters to $\eta_{s}=\eta_{n}=0.158\sqrt{mk}$ and $B=0.0001\sigma^{2}\sqrt{mk^{3}}$, respectively.
To integrate the equation of motion, we utilized the DPD-VV method [42], which is an extension of the Velocity-Verlet method to particle systems with dissipation. The time step of the simulation was set to $\Delta t=0.005\sqrt{m/k}$. The models with several different system sizes are considered: $1000\leq N\leq 512000$. For each system size, we prepared 500 different steady-state configurations for $N\leq 125000$ and 250 configurations for $N=256000,512000$. The results are all averages over these samples. Note that we used the same inherent structures for both the Stokes drag model and the normal dissipative force model.
@@ V Conclusion and future works
Our results suggest that the vibrations in granular materials are more collective than those in the thermal glass system. As mentioned in the introduction section, the vibrational dynamics of granular materials were studied experimentally, and the collective vibrations along with an increase in low-frequency modes were observed [32, 33, 34, 35]. It would be interesting to quantitatively analyze these experimental results based on those obtained in this work. However, for this purpose, our model remains too simplistic. The present study considered the normal dissipative force only, while the tangential dissipative force should also be present in granular materials. To account for this in our model, we need to treat the case where the damping matrix $\Gamma$ and the Hessian matrix $\mathcal{H}$ do not share the same matrix structure. In this case, the vibrational modes are no longer the eigenvectors of the Hessian matrix $\mathcal{H}$. Furthermore, the dynamics of real-world granular materials are affected by the friction force arising from the contacts between particles. Understanding these effects on vibrational dynamics will be the focus of future studies on granular materials.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant Numbers JP20H01868, JP20H00128, 22K03543, 23H04495, and 24H00192. In this research work, we used the “mdx: a platform for building data-empowered society”.

### v2

@@ I Introduction
The physics of the jamming transition provides a useful framework for understanding the complex dynamics of dense granular materials. Random packings of athermal, frictionless particles serve as the central model in this area of research  [19]. This model achieves mechanical stability at a specific density, known as the jamming transition [20, 21, 19, 22, 23, 24, 25, 26]. The jamming transition represents a form of critical phenomenon, where various physical properties exhibit power laws in relation to the proximity to the critical density [27, 19, 23]. Among these, the vibrational dynamics of random packings have been extensively studied. The vibrational modes of the packings are derived from the diagonalization of the dynamical matrix, which depends solely on the mass, inter-particle interactions, and the configurations of the constituent particles. In crystals, these vibrational modes manifest as plane waves, and the vibrational density of states adhere to the notable Debye law $D(\omega)\propto\omega^{d-1}$, where $d$ denotes the spatial dimension. In contrast, many vibrational modes of the random packings are spatially disordered, contributing to an excess density of states relative to the Debye law. $D(\omega)$ exhibits a plateau down to a characteristic frequency that reveals critical power-law behavior near the jamming transition. Furthermore, the spatially localized vibrational modes, referred to as quasi-localized modes, are known to emerge in the lowest frequency regime [28, 29, 30]. Importantly, the linear rheological properties of the packings are dictated by these vibrational modes [31], highlighting the significance of understanding their vibrational properties.
@@ II.2 Linearization of the equation of motion
We consider the model at the volume fraction of $\phi\simeq 0.74$ which is higher than the jamming density. Our analyses focus on such a high density system to reveal general features of the vibrations of jammed granular materials without paying too much attention to the criticality of the jamming transition. We focus on a mechanically stable configuration, or in other words, an inherent structure of the model. The position of particle $i$ in the inherent structure is denoted by ${\bm{R}}_{i}=(R_{ix},R_{iy},R_{iz})^{\top}$. Hereafter, we introduce a compact notation of a $3N$-dimensional vector like ${\bm{a}}=(\bm{a}_{1}^{\top},\bm{a}_{2}^{\top},\cdots,\bm{a}_{N}^{\top})^{\top}$. For example, the positions of particles in the inherent structure are denoted by ${\bm{R}}=(\bm{R}_{1}^{\top},\bm{R}_{2}^{\top},\cdots,\bm{R}_{N}^{\top})^{\top}$.
When the random force is sufficiently weak, the model behaves as a solid and the particles move only around their positions in the inherent structure. In this situation, the prestress (or pressure) keeps the particles in contact with each other and the interparticle contacts never be broken by random forces even if the dissipative forces are absent. Hence, we linearize the equation of motion Eq. (1) in terms of the displacements of the particles, ${\bm{u}}\equiv{\bm{r}}-{\bm{R}}$, to obtain:
@@ II.4 Simulation setup
We first obtained the inherent structure of the model in the same way as Ref. [28]. The rattler particles which have less than $4$ contacting particles were removed iteratively. Starting from the inherent structures, we performed the molecular dynamics simulations of the linearized equation of motion Eq. (7). We performed sufficiently long simulations to obtain the steady state of the system, and then performed product runs where we calculated the physical quantities defined in Sec. II.3. To calculate the kinetic energies introduced by Eq. (14), we diagonalized the hessian matrix $\mathcal{H}$ to obtain the eigenvectors and the eigenfrequencies. We set the dissipative force parameter to $\eta_{s}=\eta_{n}=0.158\sqrt{mk}$ which corresponds to typical restitution coefficient of granular particles $\varepsilon\simeq 0.7$ [42, 43]. The random force parameter was set to $B=0.0001\sigma^{2}\sqrt{mk^{3}}$ which justifies the linear approximation of the equation of motion. This corresponds to the situation in which the random forces are vanishingly weak but large enough for macroscopic motion to be clearly observed.
To integrate the equation of motion, we utilized the DPD-VV method [44], which is an extension of the Velocity-Verlet method to particle systems with dissipation. The time step of the simulation was set to $\Delta t=0.005\sqrt{m/k}$. The models with several different system sizes are considered: $1000\leq N\leq 512000$. These system sizes enable us to observe the spatially heterogeneous vibrational dynamics. For each system size, we prepared 500 different steady-state configurations for $N\leq 125000$ and 250 configurations for $N=256000,512000$. The results are all averages over these samples. Note that we used the same inherent structures for both the Stokes drag model and the normal dissipative force model.
@@ V Conclusion and future works
Our results suggest that the vibrations in granular materials are more collective than those in the thermal glass system. As mentioned in the introduction section, the vibrational dynamics of granular materials were studied experimentally, and the collective vibrations along with an increase in low-frequency modes were observed [32, 33, 34, 35]. It would be interesting to quantitatively analyze these experimental results based on those obtained in this work. However, for this purpose, our model remains too simplistic. The present study considered only the normal dissipative force, while in real experiments, there must be situations where both the normal dissipative force and Stokes drag are involved. Besides, the tangential dissipative force should also be present in granular materials. To account for this in our model, we need to treat the case where the damping matrix $\Gamma$ and the Hessian matrix $\mathcal{H}$ do not share the same matrix structure. In this case, the vibrational modes are no longer the eigenvectors of the Hessian matrix $\mathcal{H}$. Furthermore, the dynamics of real-world granular materials are affected by the friction force arising from the contacts between particles. Understanding these effects on vibrational dynamics will be the focus of future studies on granular materials.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request. Source data are provided with this paper.
Conflicts of interest
There are no conflicts of interest to declare.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant Numbers JP20H01868, JP20H00128, JP22K03543, JP23H04495, and JP24H00192. In this research work, we used the “mdx: a platform for building data-empowered society”.